A and B are the two students. Probabilities of solving a problem by A and B separately are 5/9 and 7/11 respectively. Find the probability that the problem will be solved.

Grade:Upto college level

1 Answers

Latika Leekha
9 years ago
Probability that the problem is solved by A = P (A) = 5/9
Probability that the problem is solved by B = P (B) = 7/11
Probability that the problem is not solved by A = P (AC) = 1 – 5/9 = 4/9
Probability that the problem is not solved by B = P (BC) = 1 – 7/11 = 4/11
Probability that the problem is solved = P(A) P(BC) + P(B) P(AC) + P(A)P(B)
= 5/9 . 4/11 + 7/11 . 4/9 + 5/9 . 7/11
= 83/99.
Thanks & Regards
Latika Leekha
askIITians Faculty